# Scaling LLM Planning: NL2FLOW for Parametric Problem Generation and Rigorous Evaluation

### v3

@@ Abstract
Effective agent performance relies on the ability to compose tools and agents into effective workflows. However, progress in Large Language Model (LLM) planning and reasoning is limited by the scarcity of scalable, reliable evaluation data. This study addresses this limitation by identifying a suitable workflow domain for LLM application. I introduce NL2Flow, a fully automated system for parametrically generating planning problems, which are expressed in natural language, a structured intermediate representation, and formal PDDL, and rigorously evaluating the quality of generated plans. NL2Flow generates a dataset of 2296 low-difficulty problems in automated workflow generation and evaluates multiple open-sourced, instruct-tuned LLMs without task-specific optimization or architectural modifications. Results reveal that the highest performing model achieved 86% success in generating valid plans and 69% in generating optimal plans, specifically for problems with feasible plans. Regression analysis shows that the influence of problem characteristics on plan generation is contingent on both model and prompt design. To investigate the potential of LLMs as natural language-to-JSON translators for workflow definition, and to facilitate integration with downstream symbolic computation tools and a symbolic planner, I evaluated the LLM’s translation performance on natural language workflow descriptions. I observed that translating natural language into a JSON representation of a workflow problem yielded a lower success rate than generating a plan directly, suggesting that unnecessary decomposition of the reasoning task may degrade performance and highlighting the benefit of models capable of reasoning directly from natural language to action. As LLM reasoning scales to increasingly complex problems, understanding the shifting bottlenecks and sources of error within these systems will be crucial.
@@ 1 Introduction
Automatic workflow generation is a critical capability for agentic systems. AI agents are now widely deployed to automate business processes across diverse industries, relying on tool use to access dynamic information, expand knowledge, and personalize interactions. In multi-agent systems, an orchestrating agent should compose agents to achieve goals. Effective agent performance hinges on the ability to efficiently ”select” and ”compose” tools or agents, essentially, to create effective workflows, in response to user requests. Traditional planning techniques, while conceptually sound, often fall short in addressing the scalability and adaptability requirements of these dynamic environments.
@@
Traditional AI planning systems rely on formal representations, such as those defined by Planning Domain Definition Language (PDDL). While these systems can offer guarantees like optimality, building them requires painstaking knowledge engineering to translate natural language goals into formal problem descriptions [1]. This translation process is a known bottleneck, and the complexity of maintaining and scaling these systems increases substantially with the number of tools and task complexities. For example, business automation systems can involve thousands of APIs. Consequently, symbolic planners struggle to find optimal workflows within reasonable timeframes, often requiring developers to manually reduce the scope of the problem by limiting the number of APIs considered.
Large Language Models (LLMs) present a promising alternative, demonstrating impressive capabilities in natural language understanding, generation, and reasoning. However, leveraging LLMs for reliable planning remains challenging. Existing approaches often resort to free-form plan generation, which lacks logical consistency, or rely on reinforcement learning confined to limited environments [2]. A key advancement would be the ability to reliably translate natural language planning problems into formal representations or to directly generate logically sound and executable plans (Figure 1). Successful LLM-based planning would significantly reduce the engineering effort associated with building robust natural language planning systems.
Despite their potential, evaluating and improving LLM planning capabilities is hampered by the scarcity of scalable data generation and evaluation methodologies. This study addresses this need by introducing NL2Flow [3], a scalable data generation method for workflows. NL2Flow uses random and parametric approaches to create a dataset of natural language descriptions, Python objects, and PDDL representations, eliminating the need for human or LLM annotation and enabling the creation of large-scale datasets. I demonstrate that NL2Flow can generate workflows in a domain where open-sourced, instructed-tuned LLMs without task-specific optimization or architectural modifications exhibit promising planning performance for low-difficulty problems. Establishing a domain in which large language models (LLMs) exhibit promising reasoning and planning capabilities without task-specific optimization and modifications can enable researchers to investigate the origins of their domain-dependent reasoning/planning performance.
@@ 2 Method
This study used an automated planning integration system, NL2Flow [3], to generate a dataset of workflow planning problems and evaluate the performance of large language models in generating workflows and translating workflow problems in natural language. The NL2Flow profiler module (profiler.generators.batch_data_generator) was used to programmatically generate a diverse set of planning scenarios, varying parameters. This allowed for controlled generation of data suitable for evaluating LLM performance across a range of workflow configurations. LLM inference was conducted on compute clusters equipped with NVIDIA H100, H200, and A100 GPUs (with larger models, for example, DeepSeek-V3, using up to 8 H200 GPUs with 400GB memory). LLM-generated workflows were evaluated using NL2Flow’s Debugger module (nl2flow.debug.debug.BasicDebugger), which assesses soundness, validity, and optimality.
@@ 2.1 Domain Description
Listing 1: NL2Flow domain example for booking a business trip [3]
@@
[1] ask(w3)
[2] list of papers = Author Workbench(w3, conference name)
@@
[4] invoice = Registration Bot(conference name, list of papers, username)
@@
[6] name, address, Employee ID, Passport = Travel Info Agent(w3)
[7] name, address, Employer Letter = Workday(Employee ID)
[8] visa = Visa Application(Passport, address, Employer Letter)
@@
Automated workflow completion was a core requirement for the intelligent systems investigated in this study. I focused on the NL2Flow domain, which combined service composition with goal-oriented dialogue to generate task workflows [4]. The objective was to develop a system capable of autonomously selecting and composing a sequence of actions, which are termed ‘custom’ actions, representing available services, to achieve user goals (e.g., submitting an expense report) (Listing 1). Each custom action was modeled as a parameterized function, and successful execution necessitated providing values for a defined set of input parameters. Critically, these parameter values were frequently unknown at the initiation of a workflow, representing information the system needed to acquire from the user. To address this, the system used a dialogue-based approach. When a parameter value was missing, the system generated an ‘ask’ action, a dialogue act specifically designed to elicit the necessary information. This process of identifying and requesting missing parameter values was analogous to slot-filling, a common technique in task-oriented dialogue systems, and ensured the system possessed the information required to progress. Successfully executing a custom action updated the system’s internal state, transitioning the corresponding parameter from unknown to known, and enabling the system to proceed with subsequent actions in the workflow.
@@ 2.2 Parametric Data Generation
A parametric workflow problem generator was developed to create an intermediate representation (IR) of the problem space. Implemented in Python, this generator allows for controlled variation in workflow complexity and characteristics. A dataset of 2,296 workflows was generated using a fixed random seed (20250429) to ensure reproducibility. Workflow generation was governed by six key parameters (Table 1): the number of custom actions (representing available operations), custom action arity (the number of input/output parameters defining each action’s interface), the number of goal actions (specifying the desired outcome), the proportion of slot-fillable variables (variables requiring values during planning), the complexity of variable mapping, and the proportion of coupled actions (actions where outputs serve as inputs to subsequent actions, creating dependencies). I aimed to generate ten unique samples for each valid parameter combination. Custom actions were defined as callable components with specified input and output parameter sets, invoked when all required input parameters are known, and producing corresponding output parameters. These actions represent operations commonly found in agentic frameworks, such as agents and tools.
To construct a workflow problem, 30 variables were generated and designated as either slot-fillable or initially known. Custom actions were then created and connected to establish dependencies, with action outputs serving as inputs for subsequent actions. Variables not directly used within any action were assigned pre-defined values. A custom action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the custom action set. To simplify the problem and focus on core workflow generation challenges, variable mapping was excluded from this study. The resulting intermediate representation (Python object) was compiled into PDDL strings using NL2Flow, enabling symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication.
2.2.1 Intermediate Representation
Listing 2: Intermediate representation from data generator. An example of the intermediate representation is provided in the supplementary material as ‘intermediate_representation.json‘.
available_agents”: [
agent_id”: a__2”,
actuator_signature”: {
in_sig_full”: [
name”: v__7”,
data_type”: null,
mappable”: false,
slot_fillable”: true
out_sig_full”: [
name”: v__9”,
data_type”: null,
mappable”: false,
slot_fillable”: true
goal_agent_ids”: [
a__1”,
a__0
mappings”: [],
available_data”: [
v__10”,
null
The parametric workflow problem generator constructed an intermediate representation (Listing 2), implemented as a Python object, to define the problem space. This representation comprises five key components:
  • available_agentsavailable\_agents represent the available custom actions
@@ 2.2.1 Intermediate Representation
From this intermediate representation, I generated both natural language descriptions, using predefined templates, and a formal PDDL representation of the workflow problem.
2.3 Prompt Generation
Listing 3: Verbose planning prompt

### v4

# Scaling LLM Planning: NL2FLow for Parametric Workflow Problem Generation and Rigorous Evaluation

@@ Abstract
Robust workflow composition is critical for effective agent performance, yet progress in Large Language Model (LLM) planning and reasoning is hindered by a scarcity of scalable evaluation data. This work introduces NL2Flow, a fully automated pipeline for generating and evaluating workflow planning problems. NL2Flow generates problems parametrically in a structured intermediate representation, translating them into both natural language and formal PDDL. I evaluate several open-source, instruct-tuned LLMs on a dataset of 2296 low-difficulty problems generated by NL2Flow. Results demonstrate that the best-performing model achieved 86% success in generating valid plans and 69% in generating optimal plans (for solvable problems). Regression analysis shows that the influence of problem characteristics on plan generation is contingent on both model and prompt design. Importantly, translating natural language problems into a structured JSON representation prior to symbolic planning significantly improved success rates, suggesting a benefit from neuro-symbolic integration. These findings underscore the importance of understanding error sources within LLM reasoning as systems scale to more complex tasks. As LLM reasoning scales to increasingly complex problems, understanding the shifting bottlenecks and sources of error within these systems will be crucial.
@@ 1 Introduction
Agentic systems, increasingly deployed to automate complex business processes, rely on the ability to dynamically compose tools and agents into effective workflows. The core challenge in building these systems is planning: determining the optimal sequence of actions to achieve a given goal. While traditional AI planning techniques offer formal guarantees, they struggle with the scalability and adaptability demanded by real-world applications involving a large number of tools and rapidly changing environments.
@@
Traditional AI planning systems rely on formal representations, such as those defined by Planning Domain Definition Language (PDDL). While these systems can offer guarantees like optimality, building them requires painstaking knowledge engineering to translate natural language goals into formal problem descriptions (georgievski2015htn, ). This translation process is a known bottleneck, and the complexity of maintaining and scaling these systems increases substantially with the number of tools and task complexities. For example, business automation systems can involve thousands of APIs. Consequently, symbolic planners struggle to find optimal workflows within reasonable timeframes, often requiring developers to manually reduce the scope of the problem by limiting the number of APIs considered.
Large Language Models (LLMs) present a promising alternative, demonstrating impressive capabilities in natural language understanding, generation, and reasoning. However, using LLMs for reliable planning remains challenging. Existing approaches often resort to free-form plan generation, which lacks logical consistency, or rely on reinforcement learning confined to limited environments (planningopen, ). A key advancement would be the ability to reliably translate natural language planning problems into formal representations or to directly generate logically sound and executable plans (Figure 1). Successful LLM-based planning would significantly reduce the engineering effort associated with building robust natural language planning systems.
Despite their potential, evaluating and improving LLM planning capabilities is hampered by the scarcity of scalable data generation and evaluation methodologies. This study addresses this need by introducing NL2Flow (https://github.com/IBM/nl2flow), a scalable data generation and evaluation method for workflows. NL2Flow uses random and parametric approaches to create a dataset of workflow problems, eliminating the need for human or LLM annotation and enabling the creation of large-scale datasets. I demonstrate that NL2Flow can generate workflow problems in a domain where open-sourced, instructed-tuned LLMs without task-specific optimization or architectural modifications exhibit promising planning performance for low-difficulty problems. Establishing a domain in which large language models (LLMs) exhibit promising reasoning and planning capabilities without task-specific optimization and modifications can enable researchers to investigate the causes of their domain-dependent reasoning/planning performance.
@@ 2 Method
This study used an automated planning integration system, NL2Flow, to generate a dataset of workflow planning problems and evaluate the performance of large language models in generating workflows and translating workflow problems in natural language. The NL2Flow’s profiler module (profiler.generators.batch_data_generator) was used to programmatically generate a diverse set of planning scenarios, varying parameters. This allowed for controlled generation of data suitable for evaluating LLM performance across a range of workflow configurations. LLM inference was conducted on compute clusters equipped with NVIDIA H100, H200, and A100 GPUs (with larger models, for example, DeepSeek-V3, using up to 8 H200 GPUs with 400GB memory). LLM-generated workflows were evaluated using NL2Flow’s Debugger module (nl2flow.debug.debug.BasicDebugger), which assesses soundness, validity, and optimality. A description of the AI planning principles is provided in the Appendix.
@@ 2.1 Domain Description
Listing 1: NL2Flow domain example for booking a business trip (https://github.com/IBM/nl2flow)
@@
[2] invoice = Registration Bot(conference name, username)
@@
[4] name, address, Employee ID, Passport = Travel Info Agent(username)
[5] visa = Visa Application(Passport, address)
@@
Automated workflow completion was a core requirement for the intelligent systems investigated in this study. I focused on the NL2Flow domain, which combined service composition with goal-oriented dialogue to generate task workflows (chakraborti2022natural, ). The objective was to develop a system capable of autonomously selecting and composing a sequence of actions, which are termed ‘custom’ actions, representing available services, to achieve user goals (e.g., submitting an expense report) (Listing 1). Each ‘custom‘ action was modeled as a parameterized function, and successful execution necessitated providing values for a defined set of input parameters. Critically, these parameter values were frequently unknown at the initiation of a workflow, representing information the system needed to acquire from the user. To address this, the system used a dialogue-based approach. When a parameter value was missing, the system generated an ‘ask’ action, a dialogue act specifically designed to elicit the necessary information. This process of identifying and requesting missing parameter values was analogous to slot-filling, a common technique in task-oriented dialogue systems, and ensured the system possessed the information required to progress. Successfully executing a ‘custom‘ action updated the system’s internal state, transitioning the corresponding parameter from unknown to known, and enabling the system to proceed with subsequent actions in the workflow.
@@ 2.2 Parametric Data Generation
A parametric workflow problem generator was developed to create an intermediate representation (IR) of the problem space. Implemented in Python, this generator allows for controlled variation in workflow complexity and characteristics. A dataset of 2,296 workflows was generated using a fixed random seed (20250429) to ensure reproducibility. Workflow generation was governed by six key parameters (Table 1): the number of ‘custom‘ actions (representing available operations), ‘custom‘ action arity (the number of input/output parameters defining each action’s interface), the number of goal actions (specifying the desired outcome), the proportion of slot-fillable variables (variables requiring values during planning), the complexity of variable mapping, and the proportion of coupled actions (actions where outputs serve as inputs to subsequent actions, creating dependencies). I aimed to generate ten unique samples for each valid parameter combination. ‘custom‘ actions were defined as callable components with specified input and output parameter sets, invoked when all required input parameters are known, and producing corresponding output parameters. These actions represent operations commonly found in agentic frameworks, such as agents and tools.
@@
To construct a workflow problem, 30 variables were generated and designated as either slot-fillable or initially known. ‘custom‘ actions were then created and connected to establish dependencies, with action outputs serving as inputs for subsequent actions. Variables not directly used within any action were assigned pre-defined values. A ‘custom‘ action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the ‘custom‘ action set. To focus on core workflow generation challenges, we excluded variable mapping and initial variable values from this study. The resulting intermediate representation (Python object) was compiled into PDDL strings using NL2Flow, enabling symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication.
The parametric workflow problem generator constructed an intermediate representation (Listing A1 in Appendix), implemented as a Python object, to define the problem space. This representation comprises five key components:
  • available_agentsavailable\_agents represent the available ‘custom‘ actions
@@
From this intermediate representation, I generated natural language prompts and a formal PDDL representation of the workflow problem. See Appendix, Prompt Generation and PDDL Compilation for details on prompt generation and PDDL compilation, respectively. Language model responses were parsed according to the procedure outlined in Appendix, Language Model Response Parsing.
2.3 Plan Evaluation
NL2Flow evaluated plans generated by LLMs using the Planning Domain Definition Language (PDDL). LLM-generated plans were extracted from text and converted to PDDL via NL2Flow, using an intermediate representation from the data generator. Details of this PDDL compilation process are provided in the PDDL Compilation subsection of the Appendix. The resulting PDDL formulations were then assessed with a symbolic planner. For clarity, a sound plan is defined as one that is executable; a valid plan is executable and achieves the goal state; and an optimal plan is executable, reaches the goal state, and does so with minimal cost. When a planning problem lacks a feasible solution, any proposed plan is invalid. Furthermore, identifying an optimal plan requires the existence of at least one feasible solution.
To assess the soundness of generated plans, the pddl compilation omits goal achievement. NL2Flow’s approach distinguishes between soundness and validity. For soundness checking, the planning problem was compiled into PDDL, but instead of including the original goals as requirements, the actions reported as executed within the LLM’s response were incorporated. This approach tests the LLM’s claimed actions actually executable. Validity was checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan. In this low-difficulty workflow problem setting, a plan was invalid, also, if the system prompted the user for a non-slot-fillable variable, or if a ‘custom‘ action was invoked with unknown input parameter values. I investigated the use of VAL (howey2004val, ) for plan validation; however, VAL did not support the PDDL language features necessary to represent the complexities introduced by NL2Flow’s PDDL compilation process, precluding its use in this evaluation.
[…]
Following normalization of the explanatory variables (Tables LABEL:tab:lrnoplanlong, LABEL:tab:lrnoplanshort, LABEL:tab:lrplanvaliditylong, LABEL:tab:lrplanvalidityshort, LABEL:tab:lrplanoptimallong, and LABEL:tab:lrplanoptimalshort in Appendix), I used logistic regression, implemented in the Python statsmodels package, to assess the contribution of these variables to different aspects of LLM workflow generation capacity.
3 Results
3.1 Planning Task
3.1.1 Plan Length
[…]
Kolmogorov-Smirnov tests indicated that prompt style did not significantly alter plan length distributions for DeepSeek-V3 and Llama-3.1-405b-instruct-fp8, but did significantly affect them for other models (Figures A1 and A2 in Appendix). Mann-Whitney U tests further showed that prompt style did not significantly influence the median plan length for DeepSeek-V3, Llama-3.1-405b-instruct-fp8, and Llama-3.3-70b-instruct, while it did impact the median plan length for the remaining models.
3.1.2 Sound, Valid, and Optimal Planning
3.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 3) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid plans was largely attributable to its ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of ‘custom‘ actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – showed that factors influencing the LLM’s failure to correctly identify ’no plan’ scenarios varied depending on the model and the prompt style (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort in Appendix). However, no single variable consistently predicted failure across all models or prompt styles.
3.1.4 Feasible Plan
When restricted to problems with feasible plans, Llama-3.1-405B-instruct-fp8 demonstrated the strongest performance, achieving the high rates of generating sound (91%), valid (85%), and optimal plans (69%) using the concise prompt style (Figure 4). I further investigated the influence of problem characteristics – including optimal plan length, number of ‘custom‘ actions, arity, coupling, proportion of slot-fillable variables, and number of goals – on the generation of valid plans. The relationship between these explanatory variables and plan quality (soundness, validity, and optimality) varied considerably across language models.
3.2 Natural Language to JSON Translation
DeekSeek-V3 and Llama-3.1-405B-instruct-FP8 consistently generated valid JSON strings (100%) (Figure 5). Specifically, all JSON outputs from these models included all required elements and contained no hallucinated information. In contrast, DeepSeek-coder-33B-instruct produced valid JSON reliably, exhibiting a mean of over 6 hallucinated elements and over 6 missing elements. The models Granite-3.3-8B-instruct, Llama-3.1-8B-instruct, and Qwen2.5-72B-instruct failed to generate valid JSON in over 90% of attempts.
4 Related Work
Current evaluation of reasoning capacity often focuses on final answer accuracy for benchmarks like MATH500 (lightman2023let, ), AIME24, GSM8K (cobbe2021training, ), and MATH. However, this metric overlooks the validity of the reasoning ’process’ itself; a correct answer can be achieved through flawed reasoning, effectively rewarding a successful, yet invalid, plan. Furthermore, relying solely on final answer correctness leaves the quality and structure of the reasoning process opaque. Similarly, BIRD (li2023can, ), a natural language to SQL translation benchmark, primarily relies on final answer accuracy, despite the inherent planning aspects of SQL query construction alongside elements like intent detection. GSM-Symbolic (mirzadeh2024gsm, ) attempts to address this by generating mathematical problems from symbolic templates, allowing for manual verification of solutions. ReasonAgain (yu2024reasonagain, ) uses GPT-4o to generate Python programs for GSM8K and MATH, validating reasoning by executing the generated code.
Beyond these benchmarks, several studies use “puzzle” datasets (shojaee2025illusion, ), which include Tower of Hanoi, river crossing, blocks world, and checker jumping, to assess LLM reasoning. These puzzles offer the advantage of verifiable reasoning steps through simulation and can be scaled to challenge LLMs as they saturate. However, the applicability of these abstract puzzles to real-world scenarios remains limited.
4.1 Static Datasets & Primarily Manual/LLM-as-Judge Evaluation
FlowBench (xiao2024flowbench, ) presents a static dataset of 51 scenarios across 6 domains, focusing on multi-turn interactions. Evaluation relied heavily on human verification by comparing expected and observed traces, identifying a need for improved LLM planning. Similarly, HuggingGPT (shen2023hugginggpt, ) uses a static, human-annotated dataset augmented with GPT-4 generated pseudo-labels, and uses LLM-as-a-judge for graph tasks. TaskLAMA (yuan2024tasklama, ) also provides a static dataset, created through human annotation, for structured complex task decomposition. ISG-BENCH (chen2024interleaved, ) focuses on interleaved text and image generation, using a static dataset and relying on MLLM-as-a-judge for scoring alignment, coherence, and helpfulness, grounded in human-annotated truth. WikiPlan and RecipePlan contribute static text-image pairs for task completion. These datasets, while valuable, often require significant human effort for creation and/or evaluation (lu2023multimodal, ).
4.2 Static Datasets & Programmatic Evaluation
TaskBench (shen2024taskbench, ) divides task automation into decomposition, tool selection, and parameter prediction, leveraging a static dataset of 17,331 samples generated with GPT-4. The authors used programmatic metrics (F1, Normalized Edit Distance, Rouge) for evaluation, with some partial manual inspection. This approach offers scalability due to a large sample size but can be limited by the quality and coverage of the initial dataset. Open Grounded Planning (planningopen, ) also aims to generate executable plans, providing LLMs with an extensive action library and leveraging existing datasets. Evaluation combines programmatic metrics, such as executability and pass rate, with LLM-as-a-judge for assessing completeness, feasibility, and relevance.
4.3 Algorithmically Generated Datasets & Programmatic Evaluation
WORFBENCH (qiao2024benchmarking, ) represents a step towards automated evaluation through its data generator for graph workflow structures. LLM responses are evaluated via subgraph matching. However, there is a limitation: the need to define the entire graph upfront, hindering scalability – a limitation that search-based methods, like my own, avoid by not requiring complete graph definition for data generation and evaluation.
My work extends existing research by presenting a fully automated pipeline for data generation and evaluation, overcoming the limitations of static datasets and enabling scalable, programmatic assessment of LLM workflow generation. However, the effectiveness of this approach is ultimately bounded by the capacity of the symbolic planner to evaluate workflow optimality.
5 Conclusion
This work demonstrates that current open-source, instruct-tuned LLMs can generate effective plans directly from natural language instructions, outperforming previous LLM planning studies that relied on formal PDDL representations (silver2022pddl, ; valmeekam2023planning, ; kokel2025acpbench, ). This difference likely arises from the nature of the reasoning tasks: our study focused on workflow creation, where generated actions inherently involve fixed input and output variables, contrasting with the flexible variable binding common in domains like the Blocks World. Importantly, the complexity of workflow creation is comparable to that of domains used in prior LLM planning research. The observed pattern reinforces the hypothesis that LLM planning capabilities are domain-specific, potentially rooted in their statistical reliance on token prediction and the associated sensitivity to the compositional complexity of input tokens (zhao2025chain, ).
Regression analysis confirmed that the influence of problem characteristics on the goodness of plan is dependent on both the language model and the prompting strategy employed. This highlights the importance of model- and prompt-specific analysis, cautioning against broad generalizations across LLMs. My study demonstrated that translating natural language into an intermediate JSON workflow representation significantly improved performance compared to direct plan generation (Figure 1). This suggests that prematurely replacing established neuro-symbolic planning systems with direct language-to-action reasoning may lead to a decline in plan quality. While decomposing planning into multiple steps can introduce complexity, simplifying the system by directly reasoning from language to action can, in some applications, prioritize reduced complexity over plan quality. However, as LLM reasoning capabilities advance and scale to more complex problems, the limitations and error sources within AI planning systems are likely to evolve.
NL2Flow extends beyond simple LLM assessment by providing a framework for evaluating the planning capabilities of novel agentic architectures. These architectures vary in agent and tool configurations and operate with randomly assigned goals, expressed as LLM-generated utterances based on system-defined agent and tool descriptions (including those adhering to the Model Context Protocol (MCP)). By translating these goals and configurations into PDDL, NL2Flow enables the evaluation of agentic system trajectories using a standardized planning framework, allowing for assessment of an agent’s ability to select and compose tools based on custom, symbolically-verifiable workflows. Given the prevalence of chain-of-thought, tree-of-thought, and self-consistency prompting methods in agentic systems for workflow creation, verifying the soundness of this semantic approach is crucial for improving system reliability.
Appendix
Intermediate Representation Example
Listing A1: Intermediate representation from data generator. An example of the intermediate representation is provided in the supplementary material as ‘intermediate_representation.json‘.
"available_agents": [
"agent_id": "a__2",
"actuator_signature": {
"in_sig_full": [
"name": "v__7",
"data_type": null,
"mappable": false,
"slot_fillable": true
"out_sig_full": [
"name": "v__9",
"data_type": null,
"mappable": false,
"slot_fillable": true
"goal_agent_ids": [
"a__1",
"a__0"
"mappings": [],
"available_data": [
"v__10",
null
Prompt Generation
This study examined the capacity of LLMs to generate executable plans to achieve specified goals and to translate natural language problem descriptions into JSON schemas. I adopted a few-shot prompting approach, constructing prompt styles for both tasks and providing six example demonstrations. Crucially, each example within the prompt presented a natural language description of the plan execution before the corresponding plan expressed in the target format (Listing A2). This was intended to guide the LLM by demonstrating the logical connection between action sequences and their executable representations, thereby increasing the likelihood of generating valid and goal-directed plans.
Listing A2: Planning response example
PLAN EXPLANATION
PLAN
[0] v__1 = a__2()
[1] v__2 = a__1(v__1)
[2] a__3(v__2)
For the planning task, I investigated the impact of prompt style, comparing ‘verbose‘ (Listing A5) and ‘concise‘ (Listing A6) prompt styles. The ‘verbose‘ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ‘concise‘ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. Listings A3 and A4 show the description of ‘custom‘ action a__1a\_\_1 in verbose and concise prompts, respectively. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt.
Listing A3: ‘custom‘ action in verbose prompt
Listing A4: ‘custom‘ action in concise prompt
name: a__1
inputs: v__11, v__12, v__13
outputs: v__15, v__16, v__18
For translation, a single prompt style (Listing A7) was used, designed to elicit accurate and fluent translations. The complete prompt styles used for each task – verbose_planning_prompt.txt, concise_planning_prompt.txt, and translation_prompt.txt – are included in the supplementary material to facilitate reproducibility, enable independent verification of my results, and encourage further research in this area. Each prompt style followed a consistent structure comprising three sections: Context, Instruction, and Examples (Listings A5 and A6). The Context section provided a programmatically generated natural language description of the problem. The Instruction section detailed the desired planning process, specifying preferences for plan structure, the required output format, and guidance for handling unsolvable problems. To facilitate parsing and execution, plans were to be expressed in a Python-like syntax (e.g., outputs = ⟨action name⟩(⟨parameters⟩)). I standardized action names as a__⟨number⟩and variable names as v__⟨number⟩to promote consistency and reduce ambiguity.
Listing A5: Verbose planning prompt